You wire three resistors in a circuit; a 110Ω, a 220Ω and a 440Ω across a 9.0V battery.  Determine the current in the circuit, the current in each resistor and the voltage drop in each resistor when:

  1. The resistors are wired in series.
  2. The resistors are wired in parallel.
1
Expert's answer
2021-03-16T11:36:53-0400

1) Let's first find the equivalent resistance of the resistors connected in series:


"R_{eq}=R_1+R_2+R_3,""R_{eq}=110\ \Omega+220\ \Omega+440\ \Omega=770\ \Omega."

Then, we can find the current flowing through the equivalent resistance (or the current in the circuit) from the Ohm's law:


"I=\dfrac{V}{R}=\dfrac{9\ V}{770\ \Omega}=0.012\ A."

Since the amount of current that flows through a resistors connected in series is the same, we can write:


"I=I_1=I_2=I_3=0.012\ A."

Finally, we can find the voltage drop in each resistor:


"V_1=I_1R_1=0.012\ A\cdot110\ \Omega=1.32\ V,""V_2=I_2R_2=0.012\ A\cdot220\ \Omega=2.64\ V,""V_3=I_3R_3=0.012\ A\cdot440\ \Omega=5.28\ V."

2) Let's first find the equivalent resistance of the resistors connected in parallel:


"\dfrac{1}{R_{eq}}=\dfrac{1}{R_{1}}+\dfrac{1}{R_{2}}+\dfrac{1}{R_{3}},""R_{eq}=\dfrac{1}{\dfrac{1}{R_{1}}+\dfrac{1}{R_{2}}+\dfrac{1}{R_{3}}}""R_{eq}=\dfrac{1}{\dfrac{1}{110\ \Omega}+\dfrac{1}{220\ \Omega}+\dfrac{1}{440\ \Omega}}=62.86\ \Omega."

Then, we can find the current flowing through the equivalent resistance (or the current in the circuit) from the Ohm's law:


"I=\dfrac{V}{R}=\dfrac{9\ V}{62.86\ \Omega}=0.143\ A."

Since the voltage drop is the same across all the elements in the parallel circuit, we can write:


"V=V_1=V_2=V_3=9\ V."

Finally, we can find the current in each resistor from the Ohm's law:


"I_1=\dfrac{V}{R_1}=\dfrac{9\ V}{110\ \Omega}=0.082\ A,""I_2=\dfrac{V}{R_2}=\dfrac{9\ V}{220\ \Omega}=0.041\ A,""I_3=\dfrac{V}{R_3}=\dfrac{9\ V}{440\ \Omega}=0.02\ A."